# LetsGo: Large-Scale Garage Modeling and Rendering via LiDAR-Assisted Gaussian Primitives

### v2

@@ Abstract
Figure 1: We present LetsGo - an explicit and efficient end-to-end framework for high-fidelity rendering of large-scale garages. We design a handheld Polar scanner to capture RGBD data of expansive parking environments and have scanned a garage dataset, named GarageWorld, comprising eight garages with different structures. Our LiDAR-assisted Gaussian primitives approach along with GarageWorld dataset enables various applications, such as autonomous vehicle localization, navigation and parking, as well as VFX production.
**footnotetext: Authors contributed equally to this work.footnotetext: These authors are corresponding authors.
1 Introduction
@@
However, capturing and rendering garage environments pose significant difficulties. These spaces often have low-light conditions, textureless surfaces, and a high degree of clutter (Fig.5). The intricacies of garage layouts, including narrow spaces and reflective surfaces, further complicate the modeling process. Garages with internal circular or spiral paths often lead to incomplete data and ambiguities in spatial relationships. Additionally, their extensive spatial area underscores the necessity for efficient 3D representations and lightweight rendering techniques to facilitate real-time interaction and visualization, particularly for scenarios demanding rapid situational assessment, such as navigation and path adjustment.
Existing methods for garage modeling typically rely on either manual measurements or conventional photogrammetry and LiDAR scanning. Traditional computer vision techniques, such as Structure from Motion (SfM)[71, 65] and Multi-view Stereo (MVS[91, 24]), often struggle in these environments due to the prevalence of texture-less regions and repetitive structural designs. These methods frequently fail to extract sufficient feature points and establish accurate feature correspondences necessary for estimating camera poses. Active sensing technologies based on LiDAR can calculate camera poses and scene geometry using SLAM algorithms, but the reflective materials and transparent car windows common in garages lead to geometric inaccuracies. Moreover, LiDAR data tend to be sparse, containing many holes that corrupt high-frequency textures essential for rendering the color appearance of the scene.
While recent advances in neural representation, particularly Neural Radiance Fields (NeRF)[51], have shown promise in producing high-quality renderings, they come with high computational costs and lengthy training times. Although enhancements [72, 95, 88, 16] to NeRF aim to optimize training duration and visual rendering quality, integrating such implicit representations into conventional graphics rendering pipelines and tools for rapid 3D content applications remains challenging. The emerging 3D Gaussian Splatting (3DGS)[34] method revisits explicit representations, using 3D Gaussians to articulate the geometry and appearance of scenes, achieving high-quality scene modeling and rendering. Recent works [35, 47, 61, 48, 70] extend the 3DGS approach to model large-scale outdoor scenes, achieving impressive results. However, few methods address the unique challenges of modeling large-scale indoor scenes like underground garages, where low lighting, large texture-less regions, and repetitive patterns complicate the establishment of sufficient feature correspondences between different images.
@@
During rendering, we introduce a novel level selection strategy that optimizes the trade-off between visual fidelity and device performance by considering the distance between 3D Gaussians and the rendering viewpoint. With our multi-resolution 3D Gaussian framework, we develop an LOD PC viewer that achieves rendering speeds up to four times faster than traditional 3DGS viewers on high-performance GPUs (e.g., RTX 3090). Additionally, we offer a lightweight web renderer designed to support LOD rendering across various consumer-level devices, including laptops and tablets. We will release our source codes, including our training code, high-performance PC viewer, and lightweight web viewer, to facilitate reproducible research.
Our results, gathered from the GarageWorld dataset as well as ScanNet++[13] and KITTI-360[44] datasets, indicate that our approach not only surpasses other methods in rendering quality but also maintains high rendering efficiency. The GarageWorld dataset and the LetsGo framework for large-scale garage modeling and rendering enable various applications, including autonomous driving, localization, navigation, visual effects, etc.
2 Related Work
@@
Conventional algorithms for reconstructing large-scale scenes include Structure from Motion (SfM) [82, 74, 52], Simultaneous Localization and Mapping (SLAM) [40, 4, 5, 6] and Multi-View Stereo (MVS) [68, 25, 23]. They are dedicated to discerning the three-dimensional structure of a scene through the sequential or multi-view analysis of two-dimensional image frames. All these methods leverage feature tracking and multi-view consistency to recover the 3D scene structures. SfM- and SLAM-based methods [12, 1, 19, 28, 78, 64, 42] estimate poses of input images and recover the scene structure jointly. However, their main purpose is pose estimation, and the recovered scene point clouds are always sparse, making it difficult for high-quality free-view synthesis. While MVS-based methods [41, 15, 29, 90, 87, 86], especially deep-based approaches, compute a dense depth map for each input image, the constructed scenes often lack accuracy and robustness in texture-less and complex scenes.
NeRF and 3D Gaussian Splatting Variants.
Recent advances [55, 9, 93, 36, 58, 98, 97] in neural scene representation have significantly impacted novel view synthesis. Neural Radiance Fields (NeRF), introduced by Mildenhall et al. [51], have revolutionized 3D reconstruction with a novel framework for detailed scene capture. Subsequent improvements [73, 38, 62, 60] have aimed at enhancing the visual fidelity and computational efficiency of NeRF. MipNeRF [2] and MipNeRF360 [3] address aliasing via a novel conical frustum rendering technique and a nonlinear scene representation, respectively. Techniques such as Plexnoxels [20], Instant-NGP [53], and TensoRF [8] have expedited rendering by integrating explicit encoding methods with compact MLP networks. Furthermore, Block-NeRF [75] and Mega-NeRF [79] facilitate the application of NeRF to extensive scenes. F2-NeRF [81] introduces a space-warping method for handling arbitrary camera trajectories. ScaNeRF [84] optimizes camera pose and scene representation jointly to address pose drift in large-scale scene reconstruction. Unlike these NeRF-based approaches, our method employs an explicit 3D Gaussian representation with depth priors from our RGBD scanner, enhancing realism and efficiency.
3D Gaussian Splatting (3DGS)[34] is a transformative approach characterized by its efficiency and lifelike visual quality. Extensive research [48, 54, 33, 49, 7, 83, 31, 77] has built on 3DGS. Mip-Splatting [94] mitigates artifacts from varying view sampling rates, and GaussianPro [10], which enhances Gaussian distribution with 2D image constraints. DN-Splatter [80] and DNGaussian [43] use depth and normal cues for refinement. SuGaR [27] introduces Gaussian alignment regularization for explicit mesh extraction, and GaussianSurfels [14] flattens 3D Gaussian ellipsoids into 2D ellipses for accurate surface reconstructions. 2DGS [30] transforms 3D volumes into 2D planar Gaussian disks, employing a perspective-accurate 2D splatting process aligned with geometric surfaces. Despite these advances focusing on small-scale or object-level scenes, our approach innovates with a depth regularizer and a multi-resolution Gaussian representation designed for large-scale scenes, achieving realistic visual quality and enhanced rendering efficiency.
Gaussian Splatting for Large-Scale Scenes.
To effectively manage large-scale scenes, VastGaussian [46] explores various partitioning strategies and introduces an appearance embedding module to enhance Gaussian training. Concurrent works such as DrivingGaussian [99] and Street Gaussian [89] adapt Gaussian methodologies to the dynamic contexts of urban landscapes and autonomous driving scenarios, respectively. In addition, multiple studies are integrating the Level of Detail (LOD) rendering technique with 3D Gaussian Splatting to balance rendering quality and speed. Traditional mesh-based LOD strategies [57, 11] create varying levels of detail in object meshes, selecting the appropriate level based on the viewer’s distance to the target or the desired quality. Point-based LOD methods, like Potree [66] and FastLOD [67], are designed to efficiently render large 3D point clouds, which is particularly beneficial for LiDAR data visualization.
Recently, Octree-GS [61] proposes representing Gaussians within an octree structure to render fine details at different viewing scales. CityGaussian [47] employs an LOD strategy for efficiently training and rendering of large-scale 3DGS. Hierarchical 3DGS [35] introduces a hierarchical 3DGS, showing promising results in rendering extensive walk-through datasets. However, these methods lack of on-demand rendering schemes and require high-computational devices to load complete scenes for real-time rendering, making them unsuitable for lightweight devices such as tablets or laptops. In contrast, our LetsGo method employs a multi-resolution Gaussian representation coupled with an on-demand rendering scheme. This approach achieves rendering speeds four times faster than the original 3DGS and ensures compatibility with lightweight devices, facilitating real-time rendering in large-scale garage scenes.
@@
Figure 2: Our compact Polar scanner (b) is engineered for capturing expansive garage environments (a). It is optimized for handheld operation or vehicular mounting, enabling versatile data capture in extensive spaces. At the core of Polar’s data acquisition unit lies a high-fidelity LiDAR sensor, capturing precise 3D point clouds (c), complemented by a fisheye camera that procures wide-angle 2D RGB images (d) for a complete scene modeling.
Table 1: Detailed illustrations of our GarageWorld dataset, including various datasets categorized by their environment type, geometric features, area size, number of images, point count, face count, and lighting conditions
@@
3 Garage Data Capture
3.1 Raw Data Acquisition
@@
Capturing Device. To address these problems, we design a lightweight handheld scanning device named “Polar” to jointly collect color and geometric information of the garage. The data collection unit of the Polar device comprises a color fisheye camera, a LiDAR sensor, and an IMU sensor, as visualized in Fig. 2. The fisheye camera captures RGB color information in 30 FPS. It has a resolution of 6K and a field of view (FOV) of $180\times 180$ degrees, allowing for quick and comprehensive recording of vivid color data. The LiDAR sensor collects 3D point clouds, recording geometric information at a rate of 2.6 million points per second. With a measurement accuracy of 1 to 1.5 cm and a maximum detection distance of 50 m, it is ideal for 3D scanning of large garage scenes. The IMU sensor provides acceleration information on the device’s motion, enabling more accurate pose estimation. In addition to the data acquisition unit, we also equip the Polar device with a data processing unit consisting of a mini PC for real-time SLAM calculations. The data processing unit is powered by two removable batteries that provide more than 30 minutes of single scan endurance. With this unit, one can use a smartphone to connect with the Polar device and preview 3D point cloud reconstruction results in real-time. More information about Polar device along with the calibration process is provided in Appendix.A.
@@
Scanning Scheme. We use the Polar device to scan various garages for modeling and free-viewpoint rendering. The scanning trajectory for each garage is meticulously designed to emulate drive-through and parking. We collect data along each trajectory four times: one for forward-facing, one for backward-facing, one for side-left, and one for side-right, ensuring comprehensive capturing of the garage. We set the camera to auto exposure and auto ISO to accommodate complex lighting changes in the garage. For garages that are partially open-air and partially covered, the data is collected when the sunlight is weak, i.e., during the early morning or evening. This ensures the images captured in the transition between open-air and covered areas share similar illumination. We also apply a short pause for data capturing at the transition area, allowing the sensor to adapt to different lighting conditions and thus ensuring the images are neither overexposed nor underexposed. For garages with motion-activated lights, we ensure the data is collected after the light turns on, maintaining consistent lighting conditions across all images. The travel speed for data collection is at around 1.0±0.2 m/s, with a turning speed of 15°±3 °/s. To avoid motion blur in areas with insufficient light, these parameters are reduced to 0.5 m/s and 10 °/s, respectively. Some garages are very large, for example, over 30,000 square meters. Thus, we divide the large garage into small subsections and collect data for each subsection. The data is fused later after collection.
Figure 3: Overview of our LiDAR-assisted Gaussian splatting framework. Initially, we generate a base mesh using color and depth data collected by our self-designed Polar device. The data is then partitioned into blocks for parallel and rapid processing. Next, we downsample the high-quality scanned point clouds into multi-resolution point cloud for our LOD-LiDAR-RGS (Sec.4.2) method initialization. In addition to photometric supervision, we apply our novel unbiased Gaussian depth regularizer (Sec.4.1) for geometric supervision. Finally, our system produces photorealistic LOD rendering results based on the optimized multi-resolution Gaussian representation.
3.2 GarageWorld Dataset

### v3

@@ Abstract.
Neural rendering, large-scale garage modeling, LiDAR scanning, 3D Gaussian splatting, garage dataset, level-of-detail rendering
copyright: rightsretainedjournal: TOGjournalyear: 2024journalvolume: 43journalnumber: 6publicationmonth: 12doi: 10.1145/3687762ccs: Computing methodologies Computational photographyccs: Computing methodologies Image-based rendering
Figure 1. We present LetsGo - an explicit and efficient end-to-end framework for high-fidelity rendering of large-scale garages. We design a handheld Polar scanner to capture RGBD data of expansive parking environments and have scanned a garage dataset, named GarageWorld, comprising eight garages with different structures. Our LiDAR-assisted Gaussian primitives approach along with GarageWorld dataset enables various applications, such as autonomous vehicle localization, navigation and parking, as well as VFX production.
1. Introduction
@@
However, capturing and rendering garage environments poses significant difficulties. These spaces often have low-light conditions, textureless surfaces, and a high degree of clutter (Fig. 4). The intricacies of garage layouts, including narrow spaces and reflective surfaces, further complicate the modeling process. Garages with internal circular or spiral paths often lead to incomplete data and ambiguities in spatial relationships. Additionally, their extensive spatial area underscores the necessity for efficient 3D representations and lightweight rendering techniques to facilitate real-time interaction and visualization, particularly for scenarios demanding rapid situational assessment, such as navigation and path adjustment.
Existing methods for garage modeling typically rely on either manual measurements or conventional photogrammetry and LiDAR scanning. Traditional computer vision techniques, such as Structure from Motion (SfM)  (Snavely et al., 2008; Schönberger and Frahm, 2016) and Multi-view Stereo (MVS) (Yao et al., 2018; Furukawa and Ponce, 2010), often struggle in these environments due to the prevalence of texture-less regions and repetitive structural designs. These methods frequently fail to extract sufficient feature points and establish accurate feature correspondences necessary for estimating camera poses. Active sensing technologies based on LiDAR can calculate camera poses and scene geometry using SLAM algorithms, but the reflective materials and transparent car windows common in garages lead to geometric inaccuracies. Moreover, LiDAR data tend to be sparse, containing many holes that corrupt high-frequency textures essential for rendering the color appearance of the scene.
While recent advances in neural representation, particularly Neural Radiance Fields (NeRF) (Mildenhall et al., 2021), have shown promise in producing high-quality renderings, they come with high computational costs and lengthy training times. Although enhancements (Sun et al., 2022; Zhang et al., 2020; Xu et al., 2022; Deng et al., 2022) to NeRF aim to optimize training duration and visual rendering quality, integrating such implicit representations into conventional graphics rendering pipelines and tools for rapid 3D content applications remains challenging. The emerging 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) method revisits explicit representations, using 3D Gaussians to articulate the geometry and appearance of scenes, achieving high-quality scene modeling and rendering. Recent works (Kerbl et al., 2024; Liu et al., 2024; Ren et al., 2024; Lu and Dai., 2024; Shuai et al., 2024) extend the 3DGS approach to model large-scale outdoor scenes, achieving impressive results. However, few methods address the unique challenges of modeling large-scale indoor scenes like underground garages, where low lighting, large texture-less regions, and repetitive patterns complicate the establishment of sufficient feature correspondences between different images.
@@
During rendering, we introduce a novel level selection strategy that optimizes the trade-off between visual fidelity and device performance by considering the distance between 3D Gaussians and the rendering viewpoint. With our multi-resolution 3D Gaussian framework, we develop an LOD PC viewer that achieves rendering speeds up to four times faster than traditional 3DGS viewers on high-performance GPUs (e.g., RTX 3090). Additionally, we offer a lightweight web renderer designed to support LOD rendering across various consumer-level devices, including laptops and tablets. We have released our source codes, including our training code, high-performance PC viewer, and lightweight web viewer, to facilitate reproducible research. Please refer to our project page.
Our results, gathered from the GarageWorld dataset as well as ScanNet++ (Dai et al., 2017) and KITTI-360 (Liao et al., 2022) datasets, indicate that our approach not only surpasses other methods in rendering quality but also maintains high rendering efficiency. The GarageWorld dataset and the LetsGo framework for large-scale garage modeling and rendering enable various applications, including autonomous driving, localization, navigation, visual effects, etc.
2. Related Work
@@
Conventional algorithms for reconstructing large-scale scenes include Structure from Motion (SfM) (Wu, 2013; Sweeney et al., 2016; Moulon et al., 2013), Simultaneous Localization and Mapping (SLAM) (Leonard and Durrant-Whyte, 1991; Bavle et al., 2023; Bujanca et al., 2021; Ceriani et al., 2015) and Multi-View Stereo (MVS) (Seitz et al., 2006; Goesele et al., 2007; Furukawa et al., 2010). They are dedicated to discerning the three-dimensional structure of a scene through the sequential or multi-view analysis of two-dimensional image frames. All these methods leverage feature tracking and multi-view consistency to recover the 3D scene structures. SfM- and SLAM-based methods (Crandall et al., 2011; Agarwal et al., 2011; Frahm et al., 2010; Heinly et al., 2015; Teed and Deng, 2021; Schmuck et al., 2021; Li et al., 2020) estimate poses of input images and recover the scene structure jointly. However, their main purpose is pose estimation, and the recovered scene point clouds are always sparse, making it difficult for high-quality free-view synthesis. While MVS-based methods (Li et al., 2022; Dai et al., 2019; Huang et al., 2021; Yang et al., 2021; Xu et al., 2021b, a), especially depth-based approaches, compute a dense depth map for each input image, the constructed scenes often lack accuracy and robustness in texture-less and complex scenes.
@@
Figure 2. Our compact Polar scanner (b) is engineered for capturing expansive garage environments (a). It is optimized for handheld operation or vehicular mounting, enabling versatile data capture in extensive spaces. At the core of Polar’s data acquisition unit lies a high-fidelity LiDAR sensor, capturing precise 3D point clouds (c), complemented by a fisheye camera that procures wide-angle 2D RGB images (d) for a complete scene modeling.
Table 1. Detailed illustrations of our GarageWorld dataset, including various datasets categorized by their environment type, geometric features, area size, number of images, point count, face count, and lighting conditions
@@
NeRF and 3D Gaussian Splatting Variants.
Recent advances (Noguchi et al., 2021; Chen et al., 2021; Yu et al., 2021; Kerr et al., 2023; Rebain et al., 2021; Zhao et al., 2022b, a) in neural scene representation have significantly impacted novel view synthesis. Neural Radiance Fields (NeRF), introduced by Mildenhall et al. (2021), have revolutionized 3D reconstruction with a novel framework for detailed scene capture. Subsequent improvements (Sun et al., 2023; Kulhanek and Sattler, 2023; Roessle et al., 2022; Rematas et al., 2022) have aimed at enhancing the visual fidelity and computational efficiency of NeRF. MipNeRF (Barron et al., 2021) and MipNeRF360 (Barron et al., 2022) address aliasing via a novel conical frustum rendering technique and a nonlinear scene representation, respectively. Techniques such as Plexnoxels (Fridovich-Keil et al., 2022), Instant-NGP (Müller et al., 2022), and TensoRF (Chen et al., 2022) have expedited rendering by integrating explicit encoding methods with compact MLP networks. Furthermore, Block-NeRF (Tancik et al., 2022) and Mega-NeRF (Turki et al., 2022) facilitate the application of NeRF to extensive scenes. F2-NeRF (Wang et al., 2023) introduces a space-warping method for handling arbitrary camera trajectories. ScaNeRF (Wu et al., 2023) optimizes camera pose and scene representation jointly to address pose drift in large-scale scene reconstruction. Unlike these NeRF-based approaches, our method employs an explicit 3D Gaussian representation with depth priors from our RGBD scanner, enhancing realism and efficiency.
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) is a transformative approach characterized by its efficiency and lifelike visual quality. Extensive research (Lu and Dai., 2024; Niedermayr et al., 2024; Keetha et al., 2024; Matsuki et al., 2024; Charatan et al., 2024; Wu et al., 2024; Jiang et al., 2024; Tang et al., 2023) has built on 3DGS. Mip-Splatting (Yu et al., 2024) mitigates artifacts from varying view sampling rates, and GaussianPro (Cheng et al., 2024), which enhances Gaussian distribution with 2D image constraints. DN-Splatter (Turkulainen et al., 2024) and DNGaussian (Li et al., 2024) use depth and normal cues for refinement. SuGaR (Guédon and Lepetit, 2024) introduces Gaussian alignment regularization for explicit mesh extraction, and GaussianSurfels (Dai et al., 2024) flattens 3D Gaussian ellipsoids into 2D ellipses for accurate surface reconstructions. 2DGS (Huang et al., 2024) transforms 3D volumes into 2D planar Gaussian disks, employing a perspective-accurate 2D splatting process aligned with geometric surfaces.
TRIPS (Franke et al., 2024) combines the strengths of 3D Gaussian Splatting and advanced point-based rendering (Kopanas et al., 2021; Aliev et al., 2020) by employing the concept of rasterizing points into a screen-space image pyramid, enabling high-quality reconstruction and rendering of complex details. Despite these advances focusing on small-scale or object-level scenes, our approach innovates with a depth regularizer and a multi-resolution Gaussian representation designed for large-scale scenes, achieving realistic visual quality and enhanced rendering efficiency.
Gaussian Splatting for Large-Scale Scenes.
To effectively manage large-scale scenes, VastGaussian (Lin et al., 2024) explores various partitioning strategies and introduces an appearance embedding module to enhance Gaussian training. Concurrent works such as DrivingGaussian (Zhou et al., 2024) and Street Gaussian (Yan et al., 2024) adapt Gaussian methodologies to the dynamic contexts of urban landscapes and autonomous driving scenarios, respectively. In addition, multiple studies are integrating the Level of Detail (LOD) rendering technique with 3D Gaussian Splatting to balance rendering quality and speed. Traditional mesh-based LOD strategies (Ponchio and Dellepiane, 2016; Cignoni et al., 2004) create varying levels of detail in object meshes, selecting the appropriate level based on the viewer’s distance to the target or the desired quality. Point-based LOD methods, like Potree (Schütz et al., 2016) and FastLOD (Schütz et al., 2020), are designed to efficiently render large 3D point clouds, which is particularly beneficial for LiDAR data visualization.
Recently, Octree-GS (Ren et al., 2024) proposes representing Gaussians within an octree structure to render fine details at different viewing scales. CityGaussian (Liu et al., 2024) employs an LOD strategy for efficiently training and rendering of large-scale 3DGS. However, these methods lack on-demand rendering schemes and require high-computational devices to load complete scenes for real-time rendering, making them unsuitable for lightweight devices such as tablets or laptops. Hierarchical 3DGS (Kerbl et al., 2024) delivers promising results in rendering large-scale walk-through datasets without requiring high-performance computational devices. Nonetheless, it lacks a web-based viewer that supports rendering on lightweight devices, such as laptops and iPads. In contrast, our LetsGo method employs a multi-resolution Gaussian representation coupled with an on-demand rendering scheme. This approach achieves rendering speeds four times faster than the original 3DGS and ensures compatibility with lightweight devices, facilitating real-time rendering in large-scale garage scenes.
3. Garage Data Capture
Figure 3. Overview of our LiDAR-assisted Gaussian splatting framework. Initially, we generate a base mesh using color and depth data collected by our self-designed Polar device. The data is then partitioned into blocks for parallel and rapid processing. Next, we downsample the high-quality scanned point clouds into multi-resolution point cloud for our LOD-LiDAR-RGS (Sec. 4.2) method initialization. In addition to photometric supervision, we apply our novel unbiased Gaussian depth regularizer (Sec. 4.1) for geometric supervision. Finally, our system produces photorealistic LOD rendering results based on the optimized multi-resolution Gaussian representation.
3.1. Raw Data Acquisition
@@
Capturing Device. To address these problems, we design a lightweight handheld scanning device named “Polar” to jointly collect color and geometric information of the garage. The data collection unit of the Polar device comprises a color fisheye camera, a LiDAR sensor, and an IMU sensor, as visualized in Fig. 2. The fisheye camera captures RGB color information in 30 FPS. It has a resolution of 6K and a field of view (FOV) of $180\times 180$ degrees, allowing for quick and comprehensive recording of color data. The LiDAR sensor collects 3D point clouds, recording geometric information at a rate of 2.6 million points per second. With a measurement accuracy of 1 to 1.5 cm and a maximum detection distance of 50 m, it is ideal for 3D scanning of large garage scenes. The IMU sensor provides acceleration information on the device’s motion, enabling more accurate pose estimation. In addition to the data acquisition unit, we also equip the Polar device with a data processing unit consisting of a mini PC for real-time SLAM calculations. The data processing unit is powered by two removable batteries that provide more than 30 minutes of single scan endurance. With this unit, one can use a smartphone to connect with the Polar device and preview 3D point cloud reconstruction results in real-time. More information about Polar device along with the calibration process is provided in Appendix. A.
@@
Scanning Scheme. We use the Polar device to scan various garages for modeling and free-viewpoint rendering. The scanning trajectory for each garage is meticulously designed to emulate drive-through and parking. We collect data along each trajectory four times: one for forward-facing, one for backward-facing, one for side-left, and one for side-right, ensuring comprehensive capturing of the garage. We set the camera to auto exposure and auto ISO to accommodate complex lighting changes in the garage. For garages that are partially open-air and partially covered, the data is collected when the sunlight is weak, i.e., during the early morning or evening. This ensures the images captured in the transition between open-air and covered areas share similar illumination. We also apply a short pause for data capturing at the transition area, allowing the sensor to adapt to different lighting conditions and thus ensuring the images are neither overexposed nor underexposed. For garages with motion-activated lights, we ensure the data is collected after the light turns on, maintaining consistent lighting conditions across all images. The travel speed for data collection is at around 1.0±0.2 m/s, with a turning speed of 15 °±3 °/s. To avoid motion blur in areas with insufficient light, these parameters are reduced to 0.5 m/s and 10 °/s, respectively. Some garages are very large, for example, over 30,000 square meters. Thus, we divide large garages into small subsections and collect data for each subsection. The data is fused later after collection.
3.2. GarageWorld Dataset